# Real-time Identity Defenses against Malicious Personalization of Diffusion Models

### v1

@@ Results
Fig. 4c compares the impact of different perturbation methods—Gaussian noise, RID, and optimization-based methods—on diffusion loss for both pre-trained and personalized models. Gaussian noise causes only a slight and temporary increase in diffusion loss, which diminishes during the personalization process. As a result, it has minimal effect on the quality of personalized outputs. By contrast, RID perturbations lead to a persistent and substantial increase in diffusion loss, comparable to that achieved by optimization-based methods. This effect holds consistently across LoRA-based and DB-based personalization, where RID prevents the models from effectively fine-tuning to replicate identities. These results demonstrate that RID’s learned perturbations effectively disrupt personalization while offering computational efficiency compared to optimization-based approaches.
Figure 5: RID-Ensemble can successfully defend the identities in black-box settings. a. The RID-Ensemble is trained using multiple pre-trained models to enhance its ability to defend identities across models. b. In a black-box defense scenario, personalized models are trained on defended images using previously unseen pre-trained models. The RID-Ensemble effectively protects identities even against these unknown models. c. Quantitative evaluation of RID-Ensemble’s performance shows robust identity protection in black-box defense, as measured by FID, ISM, and BRISQUE. d. The effectiveness of RID-Ensemble is further supported by increased diffusion loss across black-box models, indicating strong defense performance in preventing identity retention.
Figure 6: RID-Ensemble is robust to post-processing for defended images. a. Visual examples show that RID-Ensemble effectively defends images even after post-processing (e.g., JPEG compression and Diffpure). b. Quantitative metrics (FID, ISM, and BRISQUE) confirm that RID-Ensemble remains effective under post-processing, particularly when compared to the undefended baseline. c. Despite the impact of post-processing, RID-Ensemble maintains higher diffusion losses than undefended images, demonstrating robust defense performance.
@@
To address these challenges, we developed the RID-Ensemble. Inspired by ensemble adversarial training [34], RID-Ensemble utilizes multiple pre-trained diffusion models during training, optimizing the average of their corresponding Adv-SDS losses to ensure generalized robustness (Fig. 4a).
Black-box protecting.
Fig. 5 illustrates RID-Ensemble’s ability to defend identities in black-box scenarios. After training RID-Ensemble using multiple pre-trained models, we tested its robustness against an unseen pre-trained diffusion model, Realistic Vision (RV) [35], used for personalization.
In Fig. 5b, qualitative results demonstrate that even when adversaries fine-tune RV on RID-Ensemble-defended images, the generated outputs fail to replicate identities, showing distorted and unrecognizable features. This highlights RID-Ensemble’s generalizability to unseen models.
Quantitative evaluations in Fig. 5c further validate RID-Ensemble’s robust black-box performance. Compared to clean data, defended images achieve higher FID and BRISQUE scores and lower ISM scores, indicating stronger protection against identity retention. Importantly, RID-Ensemble consistently maintains these metrics across both seen and unseen models, demonstrating its adaptability.
Fig. 5d reveals the underlying defense mechanism. Defended images exhibit significantly higher diffusion loss on black-box models compared to clean data, and this loss remains elevated during personalization. The inability of black-box models to mitigate this loss underscores RID-Ensemble’s effectiveness in preventing identity replication, even when models outside its training set are used.
@@
Fig. 6 evaluates RID-Ensemble’s robustness to post-processing techniques that adversaries may employ to weaken protection. Two representative post-processing methods are analyzed: JPEG compression (JPEG-C) [36]: This traditional approach compresses high-frequency image information, potentially diminishing the effectiveness of added perturbations. Diffpure [37]: This diffusion-based method applies noise and then denoises defended images, leveraging the generative capacity of diffusion models to restore clean features.
@@
Fig. 6b quantitatively confirms RID-Ensemble’s post-processing robustness. After JPEG-C and Diffpure, defended images continue to achieve higher FID and BRISQUE scores and lower ISM scores compared to clean data, indicating persistent protection. Although post-processing reduces defense performance slightly, RID-Ensemble-defended images still maintain strong robustness against identity replication.
Fig. 6c analyzes diffusion losses under post-processing. While post-processing slightly decreases diffusion loss by removing some defending information, RID-Ensemble-defended images still exhibit significantly higher diffusion losses than clean data. As personalization progresses, losses for post-processed images converge to elevated levels, indicating that residual defending information remains effective in disrupting identity replication.
@@ Discussion
Diffusion models, particularly large-scale pre-trained ones, represent a major breakthrough in generative AI, rapidly transforming creative industries and gaining an ever-expanding user base. However, alongside their widespread adoption, significant security concerns have emerged. Personalization techniques, while enabling remarkable customization, can be exploited for malicious purposes such as identity fraud and unauthorized image replication. Despite the urgency of these risks, existing methods for identity defense remain impractical due to high computational demands, limiting their application in real-world scenarios.
@@
This paper introduces the Real-time Identity Defender (RID), a novel solution that addresses these challenges by enabling efficient and effective identity defense. Leveraging adversarial score distillation sampling (Adv-SDS), RID generates perturbations with a single network forward pass, providing real-time protection with minimal computational overhead. This efficiency enables local deployment on user-end devices, such as smartphones, removing the need for cloud-based processing and minimizing the risk of user data leakage. Extensive experiments demonstrate RID’s robust performance in preventing unauthorized identity replication, validating its practicality for real-world use.
@@
While RID represents a significant step forward, several challenges remain. The growing diversity of pre-trained diffusion models poses a notable challenge, as defenders trained on specific models may struggle to generalize to unseen architectures. To address this, the RID-Ensemble framework trains on multiple pre-trained models, enhancing robustness in black-box scenarios. However, RID primarily targets tuning-based personalization methods, such as Dreambooth, which involve fine-tuning model parameters. Emerging tuning-free personalization methods, such as IP-Adapter [38] and InstantID [39], introduce unique challenges that require alternative defenses. For tuning-free methods, which leverage encoders rather than fine-tuning, defenses can be applied directly to the encoder, similar to traditional adversarial methods. While tuning-based methods remain the primary focus due to their stronger personalization capabilities, integrating defenses for both paradigms is a promising direction for future research.
@@
Another critical aspect of identity defense is ensuring that defended images remain perceptually indistinguishable from clean images. While RID achieves strong protection under relatively small noise constraints, further reducing perturbation visibility is essential for practical deployment in sensitive scenarios. Future work could explore more advanced training techniques and objectives to maximize protection while adhering to stricter imperceptibility requirements. Achieving this balance is crucial for widespread adoption of identity defense technologies.
@@
Looking ahead, several research directions could further enhance the field of identity defense. One promising avenue is optimizing the search space for perturbations. Current methods operate in the pixel domain, but exploring the frequency domain could enable stronger defenses under smaller noise constraints. Additionally, defended perturbations could be transformed into meaningful signals, such as makeup patterns, achieving identity defense while simultaneously adding aesthetic effects. This shift from defensive perturbations to functional enhancements represents an exciting opportunity to expand the utility of these methods.
@@
Finally, while this paper focuses on protecting human portraits, intellectual property rights for other creative outputs, such as artwork, also warrant attention. Future research could explore methods to prevent personalization techniques from replicating the unique styles of artists without their consent. Robust defenses for artistic works would not only safeguard intellectual property but also contribute to the broader ethical use of generative AI systems.
@@
By addressing these challenges and pursuing these future directions, the field of identity defense can evolve into a practical, versatile, and ethically responsible discipline. Ensuring that generative AI systems are harnessed safely will require continued innovation, balancing technical advancements with the protection of personal and creative rights.
@@ Methods
The objective of the diffusion model is to reverse the forward process by predicting the noise added to the data at different time steps, which is formalized as follows:
@@
where $w(t)$ is a strictly positive weighting function over $t$, $\|\cdot\|_{2}$ denotes the $l_{2}$ norm, and $\epsilon_{\theta}(\cdot)$ is a noise prediction network parameterized by $\theta$. For simplicity, conditions (e.g., text prompts) are omitted in Eqn. (2). After training, samples can be generated by starting from Gaussian noise and gradually denoising through the noise prediction network.
Personalization methods [28, 29] also fine-tune pre-trained diffusion models according to the diffusion loss in Eqn. (2) (or its variants) on the customized dataset. Therefore, existing defense methods[19, 20] optimize image-specific perturbations to counteract $\mathcal{L}_{\text{Diff}}$, by solving the following optimization problem:
$\min_{\delta,\|\delta\|_{\infty}\leq\varepsilon}-\mathcal{L}_{\text{Diff}}(\theta,x+\delta),$ (3)
where $x$ is a clean image, $\delta$ is the perturbation to be optimized, $\varepsilon$ is the perturbation level (e.g., $12/255$), $\|\cdot\|_{\infty}$ is the infinity norm and $\theta$ is frozen. For instance, the update rules of AdvDM [17] and Anti-DB [19] are given by:
$\delta$ $\leftarrow\delta+\gamma\text{sign}(\nabla_{\delta}\mathcal{L}_{\text{Diff}}(\theta,x+\delta)),$ (4)
$\delta$ $\leftarrow\text{clamp}(\delta,-\varepsilon,\varepsilon),$ (5)
where $\leftarrow$ represents the assignment operation, $\gamma$ is the step size, typically set to $1/10$ of the noise constraint $\varepsilon$, and $\text{sign}(\cdot)$ denotes the sign function, which outputs values of $-1$ or $1$. The function $\text{clamp}(\cdot)$ clips the values of $\delta$ to lie within the range $[-\varepsilon,\varepsilon]$. Here, $\theta$ corresponds to the pre-trained diffusion models in AdvDM [17] and surrogate personalized models in Anti-DB [19], respectively. After optimization, the corresponding defended image for $x$ with perturbation $\delta$ is $x+\delta$.
@@
In contrast to existing defense methods, the core principle of RID lies in training a neural network to generate protective perturbations for each image via a forward inference pass, eliminating the need for individual optimization. Inspired by Differentiable Image Parameterization (DIP) [42], RID represents the defended image and its noisy versions using differentiable parameters as follows:
$x_{0}^{\text{def}}$ $=x_{0}+\delta_{\phi}(x_{0}),$ (6)
$x_{t}^{\text{def}}$ $=\alpha(t)x_{0}^{\text{def}}+\sigma(t)\epsilon,$ (7)
@@
For the training objective, a natural approach would be to attack the diffusion loss in Eqn.(2) by plugging in the defended image from Eqn.(6). Formally, using the chain rule, the gradient of $\phi$ can be expressed as:
$\nabla_{\phi}\left(-\mathcal{L}_{\text{Diff}}(\phi,x_{0})\right)=-\mathbb{E}_{t,\epsilon}[w(t)\alpha(t)(\epsilon_{\theta}(x_{t}^{\text{def}},t)-\epsilon)\frac{\partial\epsilon_{\theta}(x_{t}^{\text{def}},t)}{\partial x_{t}^{\textrm{def}}}\frac{\partial x_{0}^{\textrm{def}}}{\partial\phi}],$ (8)
where $x_{0}$ is a clean image. However, this gradient calculation involves the Jacobian $\frac{\partial\epsilon_{\theta}(x_{t}^{\text{def}},t)}{\partial x_{t}^{\textrm{def}}}$ of the pre-trained diffusion model, which is computationally intensive due to the model’s large size. Additionally, it is highly nontrivial to update $\phi$ through backpropagation if non-differentiable operations such as $\textrm{sign}(\cdot)$ and $\textrm{clamp}(\cdot)$ in Eqn. (4) on $\delta_{\phi}(x_{0})$ are introduced.
@@
We simplify the gradient in Eqn.(8) by removing the Jacobian term, resulting in the following update rule for $\phi$ given an image $x_{0}$:
$\nabla_{\phi}\mathcal{L}_{\text{Adv-SDS}}(\phi,x_{0})$ $=-\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\theta}(x_{t}^{\text{def}})-\epsilon)\frac{\partial x_{0}^{\textrm{def}}}{\partial\phi}]$ (9)
$=-\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\theta}(x_{t}^{\textrm{def}})-\epsilon)\frac{\partial\delta_{\phi}(x_{0})}{\partial\phi}],$ (10)
@@
where $q(\cdot|\cdot,t)$ is the Gaussian distribution defined by the noise-adding process in Eqn.(7) at time $t$, and $p_{\theta}(\cdot,t)$ is the corresponding marginal distribution of the diffusion model’s sampling process. Intuitively, the loss function in Eqn. (11) enforces the learned sampling process to diverge from the forward process on defended images, thereby protecting identities.
However, directly calculating the loss in Eqn. (11) is intractable because $p_{\theta}(\cdot,t)$ is parameterized by the noise prediction network. To address this, we employ a surrogate loss to monitor the optimization process and simplify the implementation:
$\mathcal{L}_{\text{Sur-Adv-SDS}}(\phi,x_{0})=\mathbb{E}_{t,\epsilon}[\left\|x_{t}^{\text{def}}-\textrm{stop-grad}(x_{t}^{\text{def}}+w(t)(\epsilon_{\theta}(x_{t}^{\text{def}},t)-\epsilon))\right\|_{2}^{2}],$ (12)
@@
Through experiments, we found that training RID solely with $\mathcal{L}_{\text{Sur-Adv-SDS}}$ achieved a reasonably high level of protection. However, the generated perturbations exhibited subtle grid-like patterns, slightly affecting the images’ visual appearance. To improve visual consistency with the original images, we introduced an additional regularization term defined as follows:
@@
Our final objective optimizes
@@
where $\lambda$ is a hyperparameter balancing the two loss functions. We set $\lambda=3$ by default to ensure both losses are of comparable magnitude. Here, $\mathcal{D}$ denotes the full training set, and $\mathcal{D}^{\prime}$ is a subset of $\mathcal{D}$ as specified later. We trained RID using automatic differentiation in PyTorch according to Eq. (14). In our experiments, the regularization term effectively mitigated the grid pattern issue. However, standalone regularization did not provide sufficient protection, as it does not leverage the priors embedded in pre-trained diffusion models directly. Further ablation studies on the loss functions are presented in the Appendix.
@@
Table 1 summarizes the settings of various RID configurations and their inference costs, measured in Gflops. As the protection performance across different RID configurations is similar, we adopt the fastest by default to minimize computational overhead.
@@
Table 1: Model architectures and computation for RID

### v2

@@ Results
Fig. 4c compares the impact of different perturbation methods, aka, Gaussian noise, RID, and optimization-based methods, on the diffusion loss for both pre-trained and personalized models. Gaussian noise causes only a slight and temporary increase in diffusion loss, which diminishes during the personalization process. As a result, it has minimal effect on the quality of personalized outputs. In contrast, RID perturbations lead to a persistent and substantial increase in diffusion loss, comparable to that achieved by optimization-based methods. This effect holds consistently across LoRA-based and DB-based personalization, where RID prevents the models from effectively fine-tuning to replicate identities. These results demonstrate that RID’s learned perturbations effectively disrupt personalization while offering computational efficiency compared to optimization-based approaches.
Figure 5: RID-Ensemble can successfully defend the identities in black-box settings. a, The RID-Ensemble is trained using multiple pre-trained models to enhance its ability to defend identities across models. b, In a black-box defense scenario, personalized models are trained on defended images using previously unseen pre-trained models. The RID-Ensemble effectively protects identities even against these unknown models. c, d, e, Quantitative evaluation of RID-Ensemble’s performance shows robust identity protection in black-box defense, as measured by FID, ISM, and BRISQUE. f, g, h, The effectiveness of RID-Ensemble is further supported by increased diffusion loss across black-box models, indicating strong defense performance in preventing identity retention. All facial images used in this figure are sourced from publicly available datasets and are permitted for academic purposes.
Figure 6: RID-Ensemble is robust to post-processing for defended images. a, Visual examples show that RID-Ensemble effectively defends images even after post-processing (e.g., JPEG compression and Diffpure). b, c, d, Quantitative metrics (FID, ISM, and BRISQUE) confirm that RID-Ensemble remains effective under post-processing, particularly when compared to the undefended baseline. e, f, g, Despite the impact of post-processing, RID-Ensemble maintains higher diffusion losses than undefended images, demonstrating robust defense performance, consistent with the results depicted in b, c, and d, respectively. All facial images used in this figure are sourced from publicly available datasets and are permitted for academic purposes.
@@
To address these challenges, we developed the RID-Ensemble inspired by ensemble adversarial training [34], which utilizes multiple pre-trained diffusion models during training, optimizing the average of their corresponding Adv-SDS losses to ensure generalized robustness (Fig. 4a).
Black-box protection.
Fig. 5 illustrates the ability of RID-Ensemble to defend identities in black-box scenarios. After training RID-Ensemble using multiple pre-trained models, we tested its robustness against an unseen pre-trained diffusion model, namely, Realistic Vision (RV) [35], used for personalization.
In Fig. 5b, the qualitative results demonstrate that even when adversaries fine-tune RV on RID-Ensemble-defended images, the generated outputs fail to replicate identities, showing distorted and unrecognizable features. This highlights RID-Ensemble’s generalizability to unseen models.
The quantitative evaluations in Fig. 5ce further validate RID-Ensemble’s robust black-box performance. Compared to clean data, defended images achieve higher FID and BRISQUE scores and lower ISM scores, indicating stronger protection against identity retention. Importantly, RID-Ensemble consistently maintains these metrics across both seen and unseen models, demonstrating its adaptability.
Fig. 5fh reveals the underlying defense mechanism. Defended images exhibit significantly higher diffusion loss on black-box models compared to clean data, and this loss remains elevated during personalization. The inability of black-box models to mitigate this loss underscores RID-Ensemble’s effectiveness in preventing identity replication, even when models outside its training set are used.
@@
Fig. 6 presents RID-Ensemble’s robustness to post-processing techniques that adversaries may employ to weaken protection. Two representative post-processing methods are analyzed: (1) JPEG compression (JPEG-C) [36], a traditional approach that compresses high-frequency image information, potentially diminishing the effectiveness of added perturbations; (2) Diffpure [37], a diffusion-based method that applies noise and then denoises defended images, leveraging the generative capacity of diffusion models to restore clean features.
@@
Fig. 6bd confirms quantitatively RID-Ensemble’s post-processing robustness. After applying JPEG-C and Diffpure, defended images continue to achieve higher FID and BRISQUE scores and lower ISM scores compared to clean data, indicating persistent protection. Although post-processing reduces defense performance slightly, RID-Ensemble-defended images still maintain strong robustness against identity replication.
Fig. 6eg analyzes diffusion losses under post-processing. While post-processing slightly decreases diffusion loss by removing some defending information, RID-Ensemble-defended images still exhibit significantly higher diffusion losses than clean data. As personalization progresses, losses for post-processed images converge to elevated levels, indicating that residual defending information remains effective in disrupting identity replication.
@@ Discussion
Diffusion models, particularly large-scale pre-trained ones, represent a major breakthrough in generative artificial intelligence (AI), rapidly transforming creative industries and gaining an ever-expanding user base. However, alongside their widespread adoption, significant security concerns have emerged. Personalization techniques, while enabling remarkable customization, can be exploited for malicious purposes such as identity fraud and unauthorized image replication. Despite the urgency of these risks, existing methods for identity defense remain impractical due to high computational demands, limiting their application in real-world scenarios.
@@
This paper introduces a novel solution called RID that addresses these challenges by enabling efficient and effective identity defense. Leveraging adversarial score distillation sampling (Adv-SDS), RID generates perturbations with a single network forward pass, providing real-time protection with minimal computational overhead. This remarkable efficiency enables local deployment on user-end devices, such as smartphones, removing the need for cloud-based processing and minimizing the risk of user data leakage. Extensive experiments demonstrate RID’s robust performance in preventing unauthorized identity replication, validating its practicality for real-world use.
@@
While RID represents a significant step forward, several challenges remain. The growing diversity of pre-trained diffusion models poses a notable challenge, as defenders trained on specific models may struggle to generalize to unseen architectures. To address this, the RID-Ensemble framework leverages multiple pre-trained models, enhancing robustness in black-box scenarios. However, RID primarily targets tuning-based personalization methods, such as Dreambooth, which involve fine-tuning model parameters. Emerging tuning-free personalization methods, such as IP-Adapter [38] and InstantID [39], introduce unique challenges that require alternative defenses. For tuning-free methods, which leverage encoders rather than fine-tuning, defenses can be applied directly to the encoder, similar to traditional adversarial methods. While tuning-based methods remain the primary focus due to their stronger personalization capabilities, integrating defenses for both paradigms is a promising direction for our future research.
@@
Another critical aspect of identity defense lies in ensuring defended images remain perceptually indistinguishable from clean images. While RID achieves promising protection under relatively small noise constraints, further reducing perturbation visibility is essential for practical deployment in sensitive scenarios. Future work could explore more advanced training techniques and objectives to maximize protection while adhering to stricter imperceptibility requirements. Achieving this balance is crucial for widespread adoption of identity defense technologies.
@@
Going forward, several research avenues hold the potential to enhance the field of identity defense. One promising avenue is optimizing the search space for perturbations. Current methods operate in the pixel domain, but exploring the frequency domain could enable stronger defenses under smaller noise constraints. Additionally, defended perturbations could be transformed into meaningful signals, such as makeup patterns, achieving identity defense while simultaneously adding aesthetic effects. This shift from defensive perturbations to functional enhancements represents an exciting opportunity to expand the utility of these methods.
@@
While our current work focuses on protecting human portraits, intellectual property rights for other creative outputs, such as artwork, also warrant attention. Future research could explore the extension of such a method to prevent personalization techniques from replicating the unique styles of artists without their consent. Robust defenses for artistic works would not only safeguard intellectual property but also contribute to the broader ethical use of generative AI systems.
@@
By addressing the aforementioned challenges and pursuing these future directions, the field of identity defense can evolve into a practical, versatile, and ethically responsible discipline. Ensuring that generative AI systems are harnessed safely will require continued innovation, balancing technical advancements with the protection of personal and creative rights.
@@ Methods
The objective of a diffusion model is to reverse the forward process by predicting the noise added to the data at different time steps, which is formalized as follows:
@@
where $w(t)$ is a strictly positive weighting function over $t$, $\|\cdot\|_{2}$ denotes the $l_{2}$ norm, and $\epsilon_{\theta}(\cdot)$ is a noise prediction network parameterized by $\theta$. For simplicity, conditions (e.g., text prompts) are omitted in Eq. (2). After training, samples can be generated by starting from Gaussian noise and gradually denoising through the noise prediction network.
Personalization methods [28, 29] also fine-tune pre-trained diffusion models according to the diffusion loss in Eq. (2) (or its variants) on the customized dataset. Therefore, existing defense methods[19, 20] optimize image-specific perturbations to counteract $\mathcal{L}_{\text{Diff}}$, by solving the following optimization problem:
$\min_{\delta,\|\delta\|_{\infty}\leq\varepsilon}-\mathcal{L}_{\text{Diff}}(\theta,x_{0}+\delta),$ (3)
where $x_{0}$ is a clean image, $\delta$ is the perturbation to be optimized, $\varepsilon$ is the perturbation level (e.g., $12/255$), $\|\cdot\|_{\infty}$ is the infinity norm and $\theta$ is frozen. For instance, the update rules of AdvDM [17] and Anti-DB [19] are given by:
$\delta$ $\leftarrow\delta+\gamma\text{sign}(\nabla_{\delta}\mathcal{L}_{\text{Diff}}(\theta,x_{0}^{\textrm{def}})),$ (4a)
$\delta$ $\leftarrow\text{clamp}(\delta,-\varepsilon,\varepsilon),$ (4b)
where $\leftarrow$ represents the assignment operation, $\gamma$ is the step size, typically set to $1/10$ of the noise constraint $\varepsilon$, $x_{0}^{\textrm{def}}\triangleq x_{0}+\delta$ is the the corresponding defended image and $\text{sign}(\cdot)$ denotes the sign function, which outputs values of $-1$ or $1$. The function $\text{clamp}(\cdot)$ clips the values of $\delta$ to lie within the range $[-\varepsilon,\varepsilon]$. Here, $\theta$ corresponds to the pre-trained diffusion models in AdvDM [17] and surrogate personalized models in Anti-DB [19], respectively.
@@
In contrast to existing defense methods, the core principle of RID lies in training a neural network to generate protective perturbations for each image via a forward inference pass, eliminating the need for individual optimization. Inspired by Differentiable Image Parameterization (DIP) [42], RID represents the defended image and its noisy versions using differentiable parameters below:
$x_{0}^{\text{def}}$ $=x_{0}+\delta_{\phi}(x_{0}),$ (5a)
$x_{t}^{\text{def}}$ $=\alpha(t)x_{0}^{\text{def}}+\sigma(t)\epsilon,$ (5b)
@@
For the training objective, a natural approach would be to attack the diffusion loss in Eq. (2) by plugging in the defended image from Eq. (5a). Formally, using the chain rule, the gradient of $\phi$ can be expressed as:
$\nabla_{\phi}\left(-\mathcal{L}_{\text{Diff}}(\phi,x_{0})\right)=-\mathbb{E}_{t,\epsilon}\left[w(t)\alpha(t)(\epsilon_{\theta}(x_{t}^{\text{def}},t)-\epsilon)\frac{\partial\epsilon_{\theta}(x_{t}^{\text{def}},t)}{\partial x_{t}^{\textrm{def}}}\frac{\partial x_{0}^{\textrm{def}}}{\partial\phi}\right],$ (6)
where $x_{0}$ is a clean image. However, this gradient calculation involves the Jacobian $\partial\epsilon_{\theta}(x_{t}^{\text{def}},t)/\partial x_{t}^{\textrm{def}}$ of the pre-trained diffusion model, which is computationally intensive due to the model’s large size. Additionally, it is highly nontrivial to update $\phi$ through backpropagation if non-differentiable operations such as $\textrm{sign}(\cdot)$ and $\textrm{clamp}(\cdot)$ in Eq. (4a) on $\delta_{\phi}(x_{0})$ are introduced.
@@
We simplify the gradient in Eq. (6) by removing the Jacobian term, resulting in the following update rule for $\phi$ given an image $x_{0}$:
$\nabla_{\phi}\mathcal{L}_{\text{Adv-SDS}}(\phi,x_{0})$ $=-\mathbb{E}_{t,\epsilon}\left[w(t)\left(\epsilon_{\theta}(x_{t}^{\text{def}})-\epsilon\right)\frac{\partial x_{0}^{\textrm{def}}}{\partial\phi}\right],$ (7)
$=-\mathbb{E}_{t,\epsilon}\left[w(t)\left(\epsilon_{\theta}(x_{t}^{\textrm{def}})-\epsilon\right)\frac{\partial\delta_{\phi}(x_{0})}{\partial\phi}\right],$
@@
where $q(\cdot|\cdot,t)$ is the Gaussian distribution defined by the noise-adding process in Eq. (5b) at time $t$, and $p_{\theta}(\cdot,t)$ is the corresponding marginal distribution of the diffusion model’s sampling process. Intuitively, the loss function in Eq. (8) enforces the learned sampling process to diverge from the forward process on defended images, thereby protecting identities.
However, directly calculating the loss in Eq. (8) is intractable because $p_{\theta}(\cdot,t)$ is parameterized by the noise prediction network. To address this, we employ a surrogate loss to monitor the optimization process and simplify the implementation:
$\mathcal{L}_{\text{Sur-Adv-SDS}}(\phi,x_{0})=\mathbb{E}_{t,\epsilon}\left[\left\|x_{t}^{\text{def}}-\textrm{stop-grad}\left(x_{t}^{\text{def}}+w(t)\left(\epsilon_{\theta}(x_{t}^{\text{def}},t)-\epsilon\right)\right)\right\|_{2}^{2}\right],$ (9)
@@
Through experiments, we found that training RID solely with $\mathcal{L}_{\text{Sur-Adv-SDS}}$ achieved a reasonably high level of protection. However, the generated perturbations exhibited subtle grid-like patterns, slightly affecting the images’ visual appearance (see Extended Data Fig. 2a). To improve visual consistency with the original images, we introduced an additional regularization term defined as:
@@
Our final objective is to optimize
@@
where $\lambda$ is a hyperparameter balancing the two loss functions. We set $\lambda=3$ by default to ensure both losses are of comparable magnitude. Here, $\mathcal{D}$ denotes the full training set, and $\mathcal{D}^{\prime}$ is a subset of $\mathcal{D}$ as specified later. We trained RID using automatic differentiation in PyTorch according to Eq. (11). In our experiments, the regularization term effectively mitigated the grid-like artifact issue. However, when applied alone, it failed to provide adequate protection—identity replication remained possible, albeit with reduced realism, as shown in Extended Data Fig. 2b. In contrast, the combined loss function delivers the best qualitative and quantitative performance to ensure robust protection, as demonstrated in Extended Data Fig. 2b-c.
@@
Extended Data Table 1 summarizes the settings of various RID configurations and their inference costs, measured in Gflops. As the protection performance across different RID configurations is similar, we adopt the fastest by default to minimize computational overhead.
@@
RID-Ensemble
$x_{0}^{\text{def-ens}}$ $=x_{0}+\delta_{\phi_{\text{ens}}}(x_{0}),$ (12a)
$x_{t}^{\text{def},(i)}$ $=\alpha_{i}(t)x_{0}^{\text{def-ens}}+\sigma_{i}(t)\epsilon,$ (12b)
To train RID-Ensemble, following Eq. (9), we optimize the surrogate loss for the $i$-th pre-trained diffusion model as:
$\mathcal{L}_{\text{Sur-Adv-SDS-}(i)}(\phi_{\text{ens}},x_{0})=\mathbb{E}_{t,\epsilon}\left[\left\|x_{t}^{\text{def-ens}}-\textrm{stop-grad}\left(x_{t}^{\text{def},(i)}+w(t)\left(\epsilon_{\theta_{(i)}}(x_{t}^{\text{def},(i)},t)-\epsilon\right)\right)\right\|_{2}^{2}\right],$ (13)
where the training datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ as well as the value of $\lambda$ are the same as in Eq. (11). In our default setting, we have $N=3$ with SD v1.4, SD v1.5 and SD v2.1.
Evaluation metrics
In this section, we present the metrics used to evaluate the protection performance.
To evaluate image quality, we utilize the widely-adopted Frechet Inception Distance (FID) metric [31], which has become a standard benchmark in image generation tasks. The FID metric operates by first extracting features using the Inception network [44], then modeling the feature distribution with a multivariate Gaussian model. The distance between the generated and reference samples is subsequently calculated based on these feature distributions as follows:
$\textrm{FID}=\|\mu_{g}-\mu_{r}\|_{2}^{2}+\textrm{tr}\left(\Sigma_{g}+\Sigma_{r}-2(\Sigma_{g}\Sigma_{r})^{1/2}\right),$ (15)
where $\mu_{g}$ and $\mu_{r}$ represent the mean feature vectors of the generated samples and reference samples, respectively, and $\Sigma_{g}$ and $\Sigma_{r}$ the covariance matrices of the features for the generated and reference samples. tr refers to the trace operator. A lower FID score indicates a smaller distance between the generated and real image distributions, corresponding to more realistic generated samples.
To evaluate the similarity between the generated images and the reference images, we employ the Identity Score Matching (ISM) metric [19]. ISM utilizes ArcFace [32], a strong recognition model, to detect and extract face embeddings from both the generated and reference images. The similarity score is computed as:
$\textrm{ISM}=\frac{\bar{\textbf{e}}_{g}\cdot\bar{\textbf{e}}_{r}}{\left\|\bar{\textbf{e}}_{g}\right\|_{2}\left\|\bar{\textbf{e}}_{r}\right\|_{2}},$ (16)
[…]
BRISQUE [33] is a reference-free image quality assessment method, where higher scores indicate lower perceptual quality. It evaluates image quality by analyzing natural scene statistics in the spatial domain. First, it computes Mean Subtracted Contrast Normalized (MSCN) coefficients:
$\hat{I}(i,j)=\frac{I(i,j)-\mu(i,j)}{\sigma(i,j)+C},$ (17)
where $I(i,j)$ is the pixel intensity, $\mu(i,j)$ and $\sigma(i,j)$ are the local mean and variance, respectively, and $C$ is a small constant to avoid division by zero. The local mean and variance are computed using a Gaussian kernel $W$ as follows
$\mu=W*I,\quad\sigma=\sqrt{W*(I-\mu)^{2}},$ (18)
where * denotes the convolution operator. Next, pairwise products of MSCN coefficients are computed in four orientations: horizontal (H), vertical (V), left diagonal (D1), and right diagonal (D2):
$H(i,j)$ $=\hat{I}(i,j)\cdot\hat{I}(i,j+1),$ (19a)
$V(i,j)$ $=\hat{I}(i,j)\cdot\hat{I}(i+1,j),$ (19b)
$D1(i,j)$ $=\hat{I}(i,j)\cdot\hat{I}(i+1,j+1),$ (19c)
$D2(i,j)$ $=\hat{I}(i,j)\cdot\hat{I}(i+1,j-1).$ (19d)
The MSCN coefficients and pairwise product images are fitted to the generalized Gaussian distribution and asymmetric generalized Gaussian distribution, respectively, to extract shape and variance parameters. These features are used to train a classifier that predicts perceptual quality. Lower BRISQUE scores indicate higher image quality. In the context of defended images, increased BRISQUE scores reflect perceptual degradation, indicating a successful defense.
Data availability
All facial data used in this study were obtained from two publicly available and widely used datasets for machine learning. The licenses associated with these datasets permit their use for non-commercial academic purposes. The datasets and licenses are accessible on GitHub at https://github.com/CelebV-HQ/CelebV-HQ and https://github.com/cydonia999/VGGFace2-pytorch.
Code availability
Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. 92470118, No. 92270118, No. 62276269), Beijing Natural Science Foundation (No. L247030, No. 1232009), and Beijing Nova Program (No. 20220484044). The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education.
Author contributions
H.G., S.N., and C.L. proposed the initial idea. H.S. and C.L. supervised the overall project, including the article structure, and experimental design. T.P. and C.D. contributed to the adversarial techniques and experimental design. H.G. performed all experiments and provided the draft writing materials. H.S. and C.L. extensively revised and finalized the manuscript, figures, and other materials. All authors participated in the editing of the manuscript.
Corresponding authors
Hao Sun (haosun@ruc.edu.cn) and Chongxuan Li (chongxuanli@ruc.edu.cn).
Extended Data Figure 1: Comparison of training pipelines: existing optimization-based methods vs. our real-Time identity defender (RID). a, Optimization-based methods individually optimize perturbations for each image via continuous gradient ascent, resulting in significant computational overhead and prolonged defense time. b, RID’s training framework employs a DiT network that learns to generate image-specific permutations. This process is guided by two key loss functions: (1) adversarial score distillation loss (Adv-SDS) incorporates pre-trained model priors to increase the diffusion robustness of defended images, and (2) regression Loss matches RID-generated perturbations with those from precomputed optimization-based methods for ten percent of data. All facial images used in this figure are sourced from publicly available datasets and are permitted for academic purposes.
Extended Data Figure 2: The ablation study of training the RID with different losses. a, We show the defended images for all loss functions. The regularization term effectively mitigates the grid-like artifacts introduced by Adv-SDS in the protected images. b, We show the generated samples from the personalization diffusion models fine-tuned on images defended by the loss functions. The combined loss achieves the best qualitative protection performance—the identity is completely obscured in the generated images. c, Quantitative comparisons of diffusion losses on defended images further confirm that the combined loss consistently delivers the strongest protection. All facial images used in this figure are sourced from publicly available datasets and are permitted for academic purposes.
Extended Data Table 1: Model architectures and training computation for RID.